# TACO-RL: Task Aware Prompt Compression Optimization with Reinforcement Learning

### v2

@@ Abstract
To address these issues, we propose a novel and efficient reinforcement learning (RL) based task-aware prompt compression method. To ensure low latency requirements, we leverage existing Transformer encoder-based token classification model while guiding the learning process with task-specific reward signals using lightweight REINFORCE algorithm. We evaluate the performance of our method on three diverse and challenging tasks including text summarization, question answering and code summarization. We demonstrate that our RL-guided compression method improves the task performance by 8% - 260% across these three scenarios over state-of-the-art compression techniques while satisfying the same compression rate and latency requirements.
@@ 1 Introduction
To address this issue, several prompt compression techniques (that attempts to reduce the context length without loosing essential information) have been introduced. These existing work can be broadly categorized into two major threads: (a) Task-aware compression models Chuang et al. (2024), that generally finetune a task-specific decoder model that leads to high inference latency and cost; and (b) Task-agnostic compression models that either removes tokens/lexical units Jiang et al. (2023); Li et al. (2023) based on their information entropy or train a supervised binary token classification model using expert compressed examples Pan et al. (2024). Therefore, the existing solutions either fail to capture the task-specific behaviours or lead to high inference cost and latency. These challenges lead to two important research questions:
  • $\bullet$
    How can we design a prompt compression model that effectively leverages both bidirectional context and task-aware reward signals for better performance (Q1)?
  • $\bullet$
    To minimize inference latency and cost, how can we train a cheaper model with proper guidance from task-specific reward signals (Q2)?
@@
To address Q2, we build our work on the foundation laid by LLMLingua-2 Pan et al. (2024) which trained a task-agnostic encoder-based transformer model in a supervised setting for binary classification of input tokens where the target compressed prompt is generated using expensive and efficient LLM model such as GPT-4 Achiam et al. (2023). We propose a novel approach, Task-Aware Prompt Compression Optimization with Reinforcement Learning (TACO-RL), to guide the finetuned generic encoder model with task-specific reward signal (to address Q1) using on-policy reinforcement learning technique. As shown in Figure 1, during the model alignment process, we generate the task output from both the original and compressed prompt, and compute the task-specific reward signal using the divergence between these two outputs. These reward signals are then used to update the base encoder model using on-policy REINFORCE algorithm Williams (1992).
To illustrate the efficacy of our proposed method against state-of-the-art (SoTA) prompt compression methods, we conducted extensive experiments on three diverse and challenging tasks on open-source benchmark datasets: (a) text summarization on MeetingBank dataset; (b) question-answering tasks on Squad dataset and (c) code summarization on CodeSearchNet dataset. The empirical results across these tasks demonstrate that our RL-guided prompt compression method can improve the task performance by 8% - 260% over LLMLingua-2 and other benchmark approaches while ensuring same compression rate and inference cost or latency. To that end, our key contributions are as follows:
  • $\bullet$
@@ 3.1 Prompt Compression
The goal of prompt compression is to reduce the context length without loosing the essential information of original prompt. To ensure low latency and inference cost, we build our work on recently introduced LLMLingua-2 Pan et al. (2024) framework that translate the problem into a binary token classification problem using an encoder-based Transformer model. Given an input prompt $X=(x_{1},x_{2},...,x_{N})$ with $N$ tokens, the compression process is defined as:
@@
where $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},...,\mathbf{h}_{N}]$, and $\mathbf{h}_{i}\in\mathbb{R}^{d}$ is the encoded representation of token $x_{i}$. For each token, a binary classification probability is computed:
@@ 3.2 REINFORCE Algorithm
4 Methodology
4.1 TACO-RL
Encoder Model and Action Sampling.
$a_{i}\sim\text{Bernoulli}(p_{i})\quad\forall i\in\{1,\ldots,n\}$ (2)
Reward Calculation.
Here, $\mathcal{M}(y_{\text{comp}},y_{\text{orig}})$ denotes the divergence metric between the output of original and compressed prompts. $r_{0}$ is a negative constant reward for out-of-range compression. $\delta(=|\mathbf{x}_{c}|-c\cdot|\mathbf{x}|)$ denotes the divergence from expected compression rate. We use two important parameters to control the learning process, a compression flexibility controller $c$ and a tolerance threshold $L$.
Compression Flexibility Controller. The compression flexibility controller $c$ is a tunable hyperparameter that represents a baseline proportion for the number of tokens to be retained in the compressed prompt relative to the original prompt. A smaller value of $c$ enforces stricter compression (fewer tokens to be retained), while a larger value allows more tokens to be preserved. This provides flexibility in controlling the trade-off between output quality and inference cost.
Tolerance Threshold. As our policy is unconstrained, the action sampling may not always satisfy the compression ratio requirements during training process. Therefore, we define tolerance threshold parameter, $L$ to control the divergence from expected number of compressed tokens. Let $\delta$ denote the divergence between the actual and expected number of compressed tokens. To ensure a smooth learning process, we allow the divergence value $\delta$ to fall within the range $[-L,L]$ and if this criteria is met, then the task-specific positive reward signal is propagated. However, to ensure that the compressed prompt is neither excessively short nor unnecessarily long, a constant negative reward $r_{0}$ is applied to penalize extreme deviation. This mechanism stabilizes the compression process and guides the model towards generating prompts of desired lengths. It should be noted that during inference, we always satisfy the exact compression rate by sampling $c\cdot|x|$ tokens with highest probability.
Policy Optimization.
where $H(\mathbf{p})$ represents Shanon’s entropy regularization Shannon (1948) and $\lambda$ balances the tradeoff between reward and exploration. The gradient of the loss function with respect to the model parameters $\theta$ is calculated using Eq. 5:
Overall Approach.
Algorithm 1 delineates the key steps of TACO-RL. We begin with a task-agnostic encoder-based Transformer model Pan et al. (2024) for token classification. In each epoch, we generate the compressed prompt for every training data from the current encoder policy $\pi_{\theta}$ and compute the output of both original and compressed prompt using GPT-3.5-turbo model. Based on these outputs, we compute the task-specific output divergence metrics and use that as a positive reinforcement if the compression ratio requirements are met, otherwise we provide a small negative reward to penalize the constraint violation. Finally, using this reward signal we update the current policy parameter $\theta$ using Eq. 5. We execute this policy optimization process for $E$ epochs to generate the final optimized task-aware compression model $\pi^{*}_{\theta}$.
Input: Training set $\mathcal{D}$, initial encoder $\pi_{\theta}$, compression rate $c$, tolerance $L$, number of epochs $E$, Metric $\mathcal{M}$
Output: Optimized encoder policy $\pi_{\theta}^{*}$
for epoch $=1$ to $E$ do
       for $P\in\mathcal{D}$ do
             $\mathbf{H}\leftarrow\pi_{\theta}(P)$
             for $w_{i}\in P$ do
                   $p_{i}\leftarrow\text{softmax}(\mathbf{W}\mathbf{h}_{i}+\mathbf{b})$
                   $a_{i}\sim\text{Bernoulli}(p_{i})$
            $P_{c}\leftarrow\{w_{i}|a_{i}=1\}$
             $y_{\text{orig}},y_{\text{comp}}\leftarrow\text{GPT}(P),\text{GPT}(P_{c})$
             $\delta\leftarrow|P_{c}|-c|P|$
             Compute $\nabla_{\theta}\mathcal{L}$ using Eq. 5
             $\theta\leftarrow\theta$ + $\nabla_{\theta}\mathcal{L}$ ;
return $\pi_{\theta}^{*}$
  // updated policy
Algorithm 1 TACO-RL
4.2 Task-specific Rewards
4.2.1 Reward Formulation
where $q$ is the question for QA tasks.
4.2.2 Summarization Tasks
For summarization, we utilize the BLEU Papineni et al. (2002) score:
$\mathcal{M}_{\text{Sum}}=\text{BLEU}(y_{\text{comp}},y_{\text{orig}})$ (6)
As a precision-based verbatim similarity metric, BLEU effectively captures n-gram overlap between the summaries of the original and compressed contexts to promote content similarity, and minimizes introduction of inaccuracies and hallucinations.
4.2.3 Question Answering Task
[…]
@@ 4.2.3 Question Answering Task
Table 1: Performance metrics for different models across various compression rates on the MeetingBank Dataset. Values in parentheses indicate deltas from the original LLMLingua-2 baseline.
@@ 5 Experiments
To evaluate the effectiveness of our proposed approach, we conducted experiments on three tasks: text summarization, question answering, and code summarization.
5.1 Datasets
We use the following datasets:
MeetingBank Hu et al. (2023) - a text summarization dataset of meeting transcripts with $\sim$44k train examples, each of maximum 512 sequence length, and 862 test examples of various lengths.
SQuAD 2.0 Rajpurkar et al. (2016) - a question answering dataset where each example consists of a context and a corresponding question. Compression for the task is challenging because it requires preservation of different parts of the context depending on the question. We use a train subset of $\sim$34k samples and a test subset of 6k samples.
CodeSearchNet Husain et al. (2020) - a code dataset where one of the tasks is to generate summaries or docstring for Python code. Since the context is not in natural language, compression for this task presents a unique challenge. For this dataset we fine-tuned the encoder on a subset of $\sim$25k examples and then tested it on a subset of 1300 examples.
@@ 5.3 Experimental Setup
In our experiments, we employed the same architecture used in LLMLingua-2, xlm-roberta-large Conneau et al. (2020) with 561M parameters, as the backbone for our prompt compression technique. We replace the lm_head with a classification head on top. The experiments were performed on a compute instance equipped with 8 NVIDIA V100 GPUs (32 GB variants). For downstream evaluation of the compressed prompts, we utilized GPT-3.5-Turbo-1103 as our target language model. To ensure deterministic outputs and consistent results across multiple runs, we set the temperature to 0. During the fine-tuning phase, we employed a learning rate of 1e-6 in conjunction with a Cosine Annealing Loshchilov and Hutter (2017) scheduler to stabilize the training process. Further details regarding the experimental setup can be found in Appendix A.
(2x)(3x)(4x)(5x)(6x)Compression Ratio00 022 2244 4466 6688 881010 1010BLEU Score4.152.822.361.881.784.742.842.261.971.647.013.812.281.451.049.578.136.274.984.2211.239.878.187.026.01Original ContextLLMLingua-2 - Py150LLMLingua-2 - CodeSearchNetLLMLinguaTACO-RL (Rouge1)TACO-RL (Bleu)
@@ 5.5 Empirical Results
We conduct experiments on the three datasets over five different compression ratios. We observe that models trained at one choice of hyper parameters (see Table 3 in the appendix) generalize well over all compression ratios.
@@
CodeSearchNet: Figure 3 shows the results for the code summarization task. Despite the poor performance of the baseline models, our approach surpasses the results obtained with original prompts. Our method shows gains ranging from 1.3x at 2x compression to 2.6x at 6x compression in BLEU scores compared to the baseline. See Table 6 in the appendix for detailed results. A key reason for this improvement is that the compressor learns to retain only the parts of the context that docstrings tend to focus on, while the original prompts lead the downstream model to produce a more general summary.
@@ 5.6 Ablation Study
We conducted an ablation study to evaluate the impact of different divergence functions on model performance. For the summarization task, we compared the following metrics: ROUGE-1, ROUGE-L, and BLEU in Table 7. For QA, we tested three reward functions: F1 score, token-wise similarity, and a combination of both in Table 5. The token-wise similarity score is calculated based on token overlaps between the question and the context sentences. See Appendix C for a detailed description. Table 2 presents the effects of different rewards on model performance at the 2x compression rate.
@@ 6 Conclusion
In this paper, we introduce TACO-RL, a novel reinforcement learning-based prompt compression method to address the computational challenges in large language models (LLMs). Our approach leverages task-specific reward signals to fine-tune a Transformer encoder-based compression model using on-policy RL, enabling effective compression while maintaining performance. Our approach performs consistently well across high compression ratios (up to 6x) without requiring re-training. We demonstrate significant performance improvements of 63%, 45%, and up to 2.6x on the Text Summarization, QA, and Code Summarization tasks respectively, compared to strong baselines. Our work offers a promising approach for optimizing prompt compression, paving the way for more efficient and performant NLP systems.
@@
Our approach to prompt compression using reinforcement learning and encoder models has few limitations. Firstly, the fine-tuning process is sensitive to the choice of reward function, base model, and task-specific prompts. Experiments with different reward metrics, such as BLEU and Rouge in Tables 2, 6, 7, show that the choice of metric can significantly impact the performance of the encoder model. Additionally, the quality of the base model, which is trained on a similar dataset, can affect the fine-tuning results.
Secondly, the complexity of the task and size of the dataset play a crucial role in the fine-tuning process. While Figure 3 shows improved performance of the fine-tuned models on a code dataset, the performance increase drops significantly if the task is switched to code-completion from code summarization (See Appendix C). Large datasets are required for the model to learn the complexities effectively, especially when using reinforcement learning.

### v3

@@ Abstract
To address these issues, we propose a novel and efficient reinforcement learning (RL) based task-aware prompt compression method. To ensure low latency requirements, we leverage existing Transformer encoder-based token classification model while guiding the learning process with task-specific reward signals using lightweight REINFORCE algorithm. We evaluate the performance of our method on three diverse and challenging tasks including text summarization, question answering and code summarization. We demonstrate that our RL-guided compression method improves the task performance by 8% - 189% across these three scenarios over state-of-the-art compression techniques while satisfying the same compression rate and latency requirements.
@@ 1 Introduction
To address this issue, several prompt compression techniques (which attempt to reduce the context length without losing essential information) have been introduced. These existing work can be broadly categorized into two major threads: (a) Task-aware compression models Chuang et al. (2024), that generally finetune a task-specific decoder model that leads to high inference latency and cost; and (b) Task-agnostic compression models that either removes tokens/lexical units Jiang et al. (2023); Li et al. (2023) based on their information entropy or train a supervised binary token classification model using expert compressed examples Pan et al. (2024). Therefore, existing solutions either fail to capture task-specific behaviors or lead to high inference cost and latency. These challenges lead to two important research questions:
  • How can we design a prompt compression model that effectively leverages bidirectional context Devlin et al. (2019) and provides low inference latency (Q1)?
  • To minimize the computational cost needed for adapting this model to a new task, how can we efficiently train a model with proper guidance from task-specific reward signals (Q2)?
@@
To address Q1, we build our work on the foundation laid by LLMLingua-2 Pan et al. (2024) which trained a task-agnostic (while being inference latency-aware) encoder-based transformer model in a supervised setting for binary classification of input tokens where the target compressed prompt is generated using an expensive and efficient LLM model such as GPT-4 Achiam et al. (2023). We propose a novel approach, Task-Aware Prompt Compression Optimization with Reinforcement Learning (TACO-RL), to guide the finetuned generic encoder model with task-specific reward signal (to address Q2) using on-policy reinforcement learning technique. As shown in Figure 1, during the model alignment process, we generate the task output from both the original and compressed prompt, and compute the task-specific reward signal using the divergence between these two outputs. These reward signals are then used to update the base encoder model using on-policy REINFORCE algorithm Williams (1992).
To illustrate the efficacy of our proposed method against state-of-the-art (SoTA) prompt compression methods, we conducted extensive experiments on three diverse and challenging tasks on open-source benchmark datasets: (a) text summarization on MeetingBank dataset; (b) question-answering tasks on Squad dataset and (c) code summarization on CodeSearchNet dataset. The empirical results across these tasks demonstrate that our RL-guided prompt compression method can improve the task performance by 8% - 189% over LLMLingua-2 and other benchmark approaches while ensuring same compression rate and inference cost or latency. To that end, our key contributions are as follows:
@@ 3.1 Prompt Compression
The goal of prompt compression is to reduce the context length without losing the essential information of the original prompt. To ensure low latency and inference cost, we build our work on recently introduced LLMLingua-2 Pan et al. (2024) framework that translates the problem into a binary token classification problem using an encoder-based Transformer model.
A key innovation in our approach is leveraging bidirectional context to make more informed compression decisions. Unlike unidirectional models, the encoder model captures contextual information from both left and right directions, enabling a more nuanced understanding of each token’s importance. This bidirectional context is crucial in determining the relevance of a token within the broader prompt context.
Given an input prompt $X=(x_{1},x_{2},...,x_{N})$ with $N$ tokens, the compression process is defined as:
@@
where $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},...,\mathbf{h}_{N}]$, and $\mathbf{h}_{i}\in\mathbb{R}^{d}$ is the encoded representation of token $x_{i}$.
The contextual representation $\mathbf{h}_{i}$ captures the token’s significance by considering its interactions with surrounding tokens through self-attention mechanisms. This means the probability of preserving or removing a token is not determined in isolation, but by its relationship to the entire prompt context.
For each token, a binary classification probability is computed:
@@ 3.2 REINFORCE Algorithm
Table 1: Performance metrics for different models across various compression rates ($\tau$) on the MeetingBank Dataset. Values in parentheses indicate deltas from the original LLMLingua-2 baseline.
4 Methodology
4.1 TACO-RL
Encoder Model and Action Sampling.
$a_{i}\sim\text{Bernoulli}(p_{i})\quad\forall i\in\{1,\ldots,n\}$ (2)
Reward Calculation.
Here, $\mathcal{M}(y_{\text{comp}},y_{\text{orig}})$ denotes the divergence metric between the output of original and compressed prompts. $r_{0}$ is a negative constant reward for out-of-range compression. $\delta(=|\mathbf{x}_{c}|-c\cdot|\mathbf{x}|)$ denotes the divergence from expected compression. We use two important parameters to control the learning process, a compression flexibility controller $c$ and a tolerance threshold $L$.
Compression Flexibility Controller. The compression flexibility controller $c$ is a tunable hyperparameter that represents a baseline proportion for the number of tokens to be retained in the compressed prompt relative to the original prompt. A smaller value of $c$ enforces stricter compression (fewer tokens to be retained), while a larger value allows more tokens to be preserved. This provides flexibility in controlling the trade-off between output quality and inference cost. Also, this parameter is only required to guide the encoder during training and is different from the compression ratio used during inference. We show that the a model trained with a single vale for $c$, generalizes to different compression ratios during training.
Tolerance Threshold. As our policy is unconstrained, the action sampling may not always satisfy the compression ratio requirements during training process. Therefore, we define tolerance threshold parameter, $L$ to control the divergence from expected number of compressed tokens. Let $\delta$ denote the divergence between the actual and expected number of compressed tokens. To ensure a smooth learning process, we allow the divergence value $\delta$ to fall within the range $[-L,L]$ and if this criteria is met, then the task-specific positive reward signal is propagated. However, to ensure that the compressed prompt is neither excessively short nor unnecessarily long, a constant negative reward $r_{0}$ is applied to penalize extreme deviation. This mechanism stabilizes the compression process and guides the model towards generating prompts of desired lengths. It should be noted that during inference, we always satisfy the exact compression rate by sampling $\tau\cdot|x|$ tokens with highest probability.
Policy Optimization.
where $H(\mathbf{p})$ represents Shannon’s entropy regularization Shannon (1948) and $\lambda$ balances the tradeoff between reward and exploration. The gradient of the loss function with respect to the model parameters $\theta$ is calculated using Eq. 5:
Overall Approach.
Algorithm 1 describes the key steps of TACO-RL. We begin with a task-agnostic encoder-based Transformer model Pan et al. (2024) for token classification. In each epoch, we generate the compressed prompt for every training data from the current encoder policy $\pi_{\theta}$ and compute the output of both original and compressed prompt using GPT-3.5-turbo model. Based on these outputs, we compute the task-specific output divergence metrics and use that as a positive reinforcement if the compression ratio requirements are met, otherwise we provide a small negative reward to penalize the constraint violation. Finally, using this reward signal we update the current policy parameter $\theta$ using Eq. 5. We execute this policy optimization process for $E$ epochs to generate the final optimized task-aware compression model $\pi^{*}_{\theta}$.
Input: Training set $\mathcal{D}$, initial encoder $\pi_{\theta}$, compression controller $c$, tolerance $L$, number of epochs $E$, Metric $\mathcal{M}$
Output: Optimized encoder policy $\pi_{\theta}^{*}$
for epoch $=1$ to $E$ do
       for $P\in\mathcal{D}$ do
             $\mathbf{H}\leftarrow\pi_{\theta}(P)$
             for $w_{i}\in P$ do
                   $p_{i}\leftarrow\text{softmax}(\mathbf{W}\mathbf{h}_{i}+\mathbf{b})$
                   $a_{i}\sim\text{Bernoulli}(p_{i})$
            $P_{c}\leftarrow\{w_{i}|a_{i}=1\}$
             $y_{\text{orig}},y_{\text{comp}}\leftarrow\text{GPT}(P),\text{GPT}(P_{c})$
             $\delta\leftarrow|P_{c}|-c|P|$
             Compute $\nabla_{\theta}\mathcal{L}$ using Eq. 5
             $\theta\leftarrow\theta$ + $\nabla_{\theta}\mathcal{L}$ ;
return $\pi_{\theta}^{*}$
  // updated policy
Algorithm 1 TACO-RL
4.2 Task-specific Rewards
4.2.1 Reward Formulation
where $q$ is the question for QA tasks.
4.2.2 Summarization Tasks
For summarization, we utilize the BLEU Papineni et al. (2002) score:
$\mathcal{M}_{\text{Sum}}=\text{BLEU}(y_{\text{comp}},y_{\text{orig}})$ (6)
As a precision-based verbatim similarity metric, BLEU effectively captures n-gram overlap between the summaries of the original and compressed contexts to promote content similarity, and minimizes introduction of inaccuracies and hallucinations. BLEU’s inclusion of a penalty for extra tokens helps prevent gaming of the reward system Skalse et al. (2022), unlike recall-based metrics such as Rouge which could be exploited by simply generating longer outputs. This design choice results in more stable training through better-calibrated rewards.
4.2.3 Question Answering Task
[…]
@@ 5 Experiments
Our experimentation spans three different domains, each presenting unique compression challenges.
5.1 Datasets and Task Diversity
MeetingBank Hu et al. (2023) is a conversational summarization dataset with $\sim$44k train examples and 862 test samples. The dataset challenges our approach by requiring compression of complex, non-linear dialogue transcripts where contextual relevance depends on nuanced speaker interactions.
SQuAD 2.0 Rajpurkar et al. (2016) presents a more sophisticated question-answering challenge. Unlike straightforward summarization, this dataset requires selective information extraction where compression must preserve only the most relevant context for a specific question. Using a representative subset of $\sim$34k training and 6k test examples, the dataset tests the model’s ability to perform context-aware compression. This task is particularly challenging because different questions may require preserving entirely different segments of the same context.
CodeSearchNet Husain et al. (2020) presents a unique non-natural language compression challenge in code documentation. From the original large-scale dataset, we curated a subset of $\sim$25k training and 1300 test examples focusing on Python code summarization. Unlike natural language, code compression demands preserving complex syntactic structures, algorithmic logic, and domain-specific semantic nuances. The challenge lies in distilling code context into meaningful summaries by identifying key functional components, understanding the overall structure and logic flow, and retaining critical variable names and comments that convey essential information for comprehensibility.
By selecting datasets spanning summarization, question-answering, and technical documentation, we comprehensively assess our prompt compression technique. Each dataset represents a distinct real-world information compression scenario: conversational summarization, targeted information extraction, and technical context distillation.
@@ 5.3 Experimental Setup
In our experiments, we employed the same architecture used in LLMLingua-2, xlm-roberta-large Conneau et al. (2020) with 561M parameters, as the backbone for our prompt compression technique. We replace the lm_head with a classification head on top. The experiments were performed on a compute instance equipped with 8 NVIDIA V100 GPUs (32 GB variants). For downstream evaluation of the compressed prompts, we utilized GPT-3.5-Turbo-1103 as our target language model. We fixed the temperature parameter at zero to guarantee reproducible outcomes and uniform results in our experiments. During the fine-tuning phase, we employed a learning rate of 1e-6 in conjunction with a Cosine Annealing Loshchilov and Hutter (2017) scheduler to stabilize the training process. Further details regarding the experimental setup can be found in Appendix A.
@@ 5.5 Empirical Results
We conduct experiments on the three datasets over five different compression ratios. We observe that models trained at one choice of hyper parameters (see Table 3 in the appendix) generalize well over all compression ratios. We also conduct statistical significance tests on MeetingBank dataset results (see Table 9) to support our performance enhancement claims.
(2x)(3x)(4x)(5x)(6x)Compression Ratio00 055 551010 10101515 15152020 20202525 25253030 30303535 3535BLEU Score18.3011.988.937.316.5419.9012.119.357.766.7923.5813.349.066.575.3129.9722.7717.5914.1512.3835.0428.8124.2321.1318.46LLMLingua-2 - CodeSearchNetLLMLingua-2 - Py150LLMLinguaTACO-RL (Rouge1)TACO-RL (Bleu)
@@
CodeSearchNet: Figure 3 shows the results for the code summarization task. Even though the base model performs poorly, leveraging our approach, the final fine-tuned model significantly outperforms all other baselines. Our method shows gains ranging from 0.91x (91%) at 2x compression to 1.89x at 5x compression in BLEU scores compared to the LLMLingua-2 baseline. See Table 6 in the appendix for detailed results. Additionally, Appendix E highlights the differences in the compression mechanisms between TACO-RL and LLMLingua-2.
@@ 5.6 Ablation Study
We performed an ablation study to assess the impact of different reward metrics on model performance. For the text summarization task, we compared the following metrics: ROUGE-1, ROUGE-L, and BLEU, as shown in Table 2. For a detailed comparison over various compression ratios, see Table 7. For the QA task, we evaluated three reward functions: F1 score, token-wise similarity, and a combination of both, as shown in Table 2 (with additional details in Table 5). The token-wise similarity score is based on token overlaps between the question and context sentences, with a detailed explanation provided in Appendix C. Additionally, we investigate the effect of the entropy regularization term $H(\mathbf{p})$ on the fine-tuned model’s performance on the MeetingBank dataset, as detailed in Table 9.
@@ 6 Conclusion
In this paper, we introduce TACO-RL, a novel reinforcement learning-based prompt compression method to address computational challenges in large language models (LLMs). Our approach leverages task-specific reward signals to fine-tune a Transformer encoder-based compression model using on-policy RL, enabling effective compression while maintaining performance. Our approach performs consistently well across high compression ratios (up to 6x) without requiring re-training. We demonstrate significant performance improvements of 45%, 63%, and up to 1.89x on the Text Summarization, QA, and Code Summarization tasks respectively, compared to strong baselines. Our work offers a promising approach to optimizing prompt compression, paving the way for more efficient and performant NLP systems.
@@
Our approach to prompt compression using reinforcement learning and encoder models has few limitations. Firstly, the fine-tuning process is sensitive to the choice of reward function, base model, and task-specific prompts. Experiments with different reward metrics, such as BLEU and Rouge in Tables 2, 6, 7, show that the choice of metric can significantly impact the performance of the encoder model. Additionally, the quality of the base model, which is trained on a similar dataset, can significantly affect the fine-tuning results.
Secondly, the intricacies of task complexity and dataset magnitude significantly influence the fine-tuning methodology. As illustrated in Figure 3, fine-tuned models demonstrate enhanced performance on code datasets, yet this improvement diminishes markedly when transitioning from code summarization to code-completion tasks (See Appendix C). Extensive documentation strings within code contexts can potentially skew the compression dynamics, as models might develop strategies to retain doc-string tokens as a form of reward manipulation. Moreover, comprehensive and expansive datasets are essential for effectively capturing and learning the nuanced complexities, particularly in reinforcement learning scenarios.